OpenStudy (anonymous):

How to find the square root of a number? I really have a tough time solving Pythagorean Theorem questions because i can't square root a number at the end.. Is there a good and easy way to square root 1-3 digit numbers?

OpenStudy (fibonaccichick666):

so the best way I know is to break it down into numbers you know the square of, if you can

OpenStudy (primeralph):

A calculator.

OpenStudy (fibonaccichick666):

ie \[\sqrt(396)=\sqrt(2*198)=\sqrt(4*99)\]

OpenStudy (fibonaccichick666):

\[=\sqrt(4*3*33)=\sqrt(4*3*3*11)=2*3 * \sqrt(11)=6\sqrt(11)\]

OpenStudy (fibonaccichick666):

This is also breaking it down into it's prime decomposition

OpenStudy (fibonaccichick666):

as a side note

OpenStudy (fibonaccichick666):

Make sense?

OpenStudy (anonymous):

So if it were to ask me to find the square root of lets say 150, i would cut 150 in half which is 75? so Square root of 150 is 75 * 2

OpenStudy (primeralph):

,,,,prime.....

OpenStudy (anonymous):

Or would i go even further?

OpenStudy (anonymous):

Until i get to a single or two digit number

OpenStudy (primeralph):

Not really.

OpenStudy (primeralph):

Until you get only prime numbers.

OpenStudy (whpalmer4):

Are you interested in getting a decimal number, or a simplified expression? For a decimal number, use a calculator, or google, or wolframalpha.com or any number of websites. For a simplified expression, factor the quantity under the radical sign into its prime factors: 2,3,5,7,11,13,17,19,23,29,31,37,41,43, etc. 25=5*5 100 = 2*50=2*2*25=2*2*5*5 51 = 3*17 32=2*16=2*2*8=2*2*2*4=2*2*2*2*2 Now to simplify, remember that each pair of identical factors under the radical sign can be moved outside, and replaced by just one: \[\sqrt{150} = \sqrt{2*75} = \sqrt{2*3*25} = \sqrt{2*3*5*5} = 5\sqrt{2*3}=5\sqrt{6}\]

OpenStudy (whpalmer4):

\[\sqrt{16} = \sqrt{2*2*2*2} = 2*2 = 4\]\[\sqrt{144} = \sqrt{2*72} = \sqrt{2*2*36} = \sqrt{2*2*2*18} \]\[=\sqrt{2*2*2*2*9} =2*2\sqrt{3*3} = 2*2*3 = 12\]

OpenStudy (fibonaccichick666):

you want to go until you get a prime number

OpenStudy (phi):

If you want more ammunition for tackling these problems, then you should 1) learn some of the divisibility rules. (This helps you to find factors) http://www.mathsisfun.com/divisibility-rules.html 2) learn as many of the primes as you can up through 100 http://www.factmonster.com/math/numbers/prime.html 3) Learn as many of the perfect squares as you can http://www.factmonster.com/ipka/A0875883.html and of course, know your multiplication table.

OpenStudy (anonymous):

I am trying to find the length of the hypotenuse 

OpenStudy (anonymous):

But they want the square root of 169 which is hard for me to give an answer.

OpenStudy (fibonaccichick666):

so, that happens to be a special square

OpenStudy (fibonaccichick666):

one that you have to memorize

OpenStudy (phi):

169 is a number you might want to memorize.

OpenStudy (fibonaccichick666):

try calculating the squares of the prime numbers like this 4,9,16,25... Does this come up?

OpenStudy (phi):

your picture shows 15... do you mean 5 ? the triplet 3,4,5 is a right triangle almost as useful, is the 5,12,13 right triangle.

OpenStudy (anonymous):

Yes sorry i meant 5

OpenStudy (anonymous):

When they tell me to find the square root of 169 i usually start from the tens place and would multiply 11 * (1-12) then 12 * (1-12) and so forth

OpenStudy (anonymous):

So i was wondering if there was a method of finding the square root easier.

OpenStudy (phi):

The easiest way is to be familiar with the squares. you know 1*1, 2*2, and so on (hopefully up to 12*12= 144) but learn more! 13*13=169 notice that 13 is a prime number, so it would take awhile to figure out 13*13 is the answer...

OpenStudy (fibonaccichick666):

I don't have a way other than memorization for 169, sorry :/

OpenStudy (anonymous):

Are you saying the square root is always a prime number?

OpenStudy (phi):

if you did not know 144 was 12*12 you could use the divisibility rules to break it into smaller numbers. for example: it ends in an even number, so 2 divides into. Even better, the last 2 digits (44) is divisible by 4, so the whole number is divisible by 4 to get 4* 36 now take the square root of 4 and of 36 to get 2*6=12 (the square root of 144) notice that 169 can not be divided by anything (except 13) so we can't do that trick)

OpenStudy (fibonaccichick666):

no, but 169 happens to be, and the most efficient way when a number is not prime to get a simplified root is to break that number down into its prime decomposition. (the prime numbers multiplied together to get that number)

OpenStudy (anonymous):

Ok well i guess it all makes more sense now, thank you both for your quick and great answers, i wish i could give you both medals haha..